OpenStudy (anonymous):

Fan and Medal Sam has 15 coins, all nickels and dimes, with a total value of $0.90. How many of each coin does he have? (Let d = the number of dimes and n = the number of nickels). A. Write a system of equations to model the problem. (5 points) B. Solve the system. (5 points) C. State your answer with the proper label or labels. (5 points)

OpenStudy (anonymous):

I figured it out, and he has 12 Nickels, and 3 Dimes.

OpenStudy (anonymous):

d + n = 15 so n = 15- d 0.1 d + .05 n = 0.90 0.1 d + .05 (15-d) = .90 0.1d + .75 - .05 d = .90 0.1 d - .5 d = .15 .05 d = .15 d = 3 d+ n = 15 3 + n = 15 n = 12

OpenStudy (anonymous):

A. (d*x)+(n*y)=.90

OpenStudy (mathmale):

@greenlegodude57: Thank you for your input. Next time, however, please help Brian find the answer(s) HIMSELF; do NOT give away the answers or more than a fraction of the steps along the way. Brian would benefit much more from learning to solve this problem himself than from having the answer dropped into his lap.

OpenStudy (mathmale):

Brian: would you mind reading through GG's solution and then trying to solve the problem yourself, without help? You'll need these skills later!